OpenStudy (anonymous):

simplify -6[4-(x-2)]

OpenStudy (anonymous):

-6[4-(x-2)] =-24+6(x-2) =-24+6x-12 =6x-36

OpenStudy (anonymous):

-6(4-(x-2) =-6(4-x+4) -24+6x-24 =6x+48.

OpenStudy (anonymous):

sorry I forgot!

OpenStudy (anonymous):

-6(4-(x-2) =-6(4-x+2) ..............check this step of yours. -24+6x-12 =6x-36.

OpenStudy (anonymous):

i have another question

OpenStudy (anonymous):

??

OpenStudy (anonymous):

none of those answers are on my worksheet